OpenStudy (sissyedgar):

Math People, used 2 b my fave Subject... Now i practically Hate it! :P Which expression is equivalent to the given expression? 13(19 – x) A. 13 • 19 – 13x B. 1319 – 13x C. 13 • 19 – x D. 13 + 19 – 13 + x

OpenStudy (sissyedgar):

@misty1212

OpenStudy (misty1212):

lol it is not so bad dear one

OpenStudy (misty1212):

first one is \[13(19-x)\] and the distributive law says \[a(b+c)=ab+ac\] so \[13(19-x)=13\times 19-13\times x\]

OpenStudy (sissyedgar):

a.k.a...... A.

OpenStudy (sissyedgar):

Combine like terms. 8 + 6x – 11x + 12 A. 20 – 5x B. 20 – 17x\ C. 20 + 5x D. 14x + 1x

OpenStudy (sissyedgar):

uhhhg... all these ????????????

OpenStudy (misty1212):

hold on i got the question marks of death here

OpenStudy (misty1212):

ah c'mon don't be a sissy!\[8+6x-11x+12\] what is \(8+12\)

OpenStudy (sissyedgar):

ikr

OpenStudy (anonymous):

combine like terms for this one \[8+12=20\] and \[6-11=-5\]

OpenStudy (anonymous):

so final answer \[20-5x\] hope that is a choice

OpenStudy (sissyedgar):

it is thx!

OpenStudy (misty1212):

@sissyedgar you got that ? it is not really that hard, just adding is all

OpenStudy (sissyedgar):

lol.. yea but them fractions and lettes are just uhhg

OpenStudy (misty1212):

i didn't see a single fraction ...

OpenStudy (sissyedgar):

o yea... i was lukin @ the next problem! :P

OpenStudy (sissyedgar):

that fraction! :P

OpenStudy (misty1212):

\[\frac{x+2}{3}=-4\] right?

OpenStudy (sissyedgar):

rite... how u do tht tho?

OpenStudy (misty1212):

get rid of the annoying fraction by multiplying both sides by 3

OpenStudy (misty1212):

the 3 cancels on the left, on the right you get \(-4\times 3=-12\)

OpenStudy (sissyedgar):

lol...

OpenStudy (sissyedgar):

rite

OpenStudy (misty1212):

in other words the whole thing is the same as \[x+2=-12\] and you can solve that in one step

OpenStudy (sissyedgar):

x = -14?

OpenStudy (misty1212):

yup

OpenStudy (sissyedgar):

yay!

OpenStudy (sissyedgar):

Evaluate the expression by using the given values of the variables. y ÷ 3 • 2 – 8 ÷ x + 5 where x = 4 and y = 12 A. 11 B. 9 C. 21 D. 5

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (misty1212):

got a calculator?

OpenStudy (misty1212):

\[12\div 3\times 2-8\div 4+5\]

OpenStudy (misty1212):

they are trying to trick you here

OpenStudy (sissyedgar):

i c

OpenStudy (misty1212):

start with \(12\div 3\)

OpenStudy (sissyedgar):

4

OpenStudy (misty1212):

ok then \(4\times 2\)

OpenStudy (sissyedgar):

8

OpenStudy (misty1212):

ok then put that aside and compute \(8\div 4\)

OpenStudy (sissyedgar):

2

OpenStudy (misty1212):

now we knock it out \[8-2+5=?\]

OpenStudy (sissyedgar):

i add 1st rite? cause PEMDAS

OpenStudy (misty1212):

oh no pemdas is a trick bad math teacher play on unsuspecting students work form left to right

OpenStudy (sissyedgar):

o, ok 11

OpenStudy (misty1212):

\[8-2=6\\ 6+5=11\]

OpenStudy (misty1212):

yeah 11

OpenStudy (sissyedgar):



OpenStudy (misty1212):

\[\frac{z-5}{7}=-1\\ z-5=-7\] you can finish it i am sure

OpenStudy (sissyedgar):

-2

OpenStudy (sissyedgar):

ur a life savor! :P

OpenStudy (sissyedgar):

Solve. 18 = – 2m + 6 – m m=